# ALPCAHUS: Subspace Clustering for Heteroscedastic Data

### v2

@@ V-A2 Clustering Error
Table 2(h) shows the clustering quality of these algorithms for synthetic heteroscedastic data. Fig. 2 complements Table 2(h) by exploring the complete heteroscedastic landscape with a visual representation that is easier to understand. We define a method called “noisy oracle” where an oracle uses the true cluster assignments to apply PCA on the low noise data only per each cluster. Using these subspace basis estimates of each cluster, the oracle performs cluster assignments using (17). This oracle reference provides a kind of lower bound on realistic clustering performance since it can approximate subspace bases separately given the true labeling.
ALPCAHUS ($B=1$) with TIPS initialization achieved lower clustering error than KSS with TIPS initialization. For the ensemble methods ($B>1$), both EKSS and ALPCAHUS significantly improved. However, EKSS still had higher clustering error in more heteroscedastic regions. Meanwhile, ALPCAHUS remained very close to the noisy oracle, indicating it more accurately estimated the underlying true labeling. Compared to other methods like TSC and ADSSC, ALPCAHUS generally outperformed them with one exception for ADSSC ($N_{2}=1,\nu_{2}/\nu_{1}=300$). Here, ADSSC performed better than ALPCAHUS when the number of noisy points is roughly equal to the number of good points. Overall, the ensemble version of ALPCAHUS was generally more robust against heteroscedasticity compared to other subspace clustering algorithms. To summarize, the effects of data quality and data quantity on the heteroscedastic subspace clustering estimates is explored in different situations. To our knowledge, this is the first systematic analysis of heteroscedasticity effects on subspace clustering quality in the literature.
@@ V-A3 Effects of Good Data
Additionally, the effects of good data quantity on the subspace clustering problem is explored to understand at what point it becomes advantageous to use ALPCAHUS over KSS methods. The following parameters are fixed $\nu_{1}=0.1,N_{2}=500$ and the following vary $N_{1},\nu_{2}$. The reporting metric is the percentage difference of EKSS clustering error (%) - ALPCAHUS clustering error (%), i.e., $\text{error}_{\text{EKSS}}-\text{error}_{\text{ALPCAHUS}}$, meaning higher values indicate our method is better. Figure 3 shows that ALPCAHUS performed better than EKSS even up to $N_{1}=160$ which represents about 33% of the total data. ALPCAHUS never performed worse than EKSS but the advantage gap narrowed as $N_{1}$ increased. Thus, there is a wide range of conditions for which ALPCAHUS is preferable to homoscedastic methods.
@@ V-A4 Clustering Initialization
Section IV proposed using the TIPS initialization scheme in this heteroscedastic context since the dot product metric used in constructing the affinity matrix in (29) is provably robust to heteroscedastic noise. The following parameters were fixed $N_{1},N_{2}$ and the following varied $\nu_{2}$ while keeping $\nu_{1}=0.1$. The noisy oracle result is included to establish a realistic performance baseline. Figure 4 shows that TIPS initialization ALPCAHUS $(B=1)$ outperformed random initialization for the non-ensemble ALPCAHUS method $(B=1)$ across the entire heteroscedastic landscape. Thus, TIPS should always be used for the non-ensemble methods for clustering quality. The ensemble version of ALPCAHUS $(B>1)$ is not included in these results because the ensemble process inherently takes advantage of random initialization to achieve a different labeling each trial.
@@ V-A5 Rank Estimation of Clusters
In Sec. IV, it is proposed to use random sign flipping to adaptively find and shrink the subspace dimension of clusters when subspace dimension is unknown. For Fig. 5 synthetic subspaces are generated with true rank $d=6$ to explore how the initial rank parameter affects the ability to estimate the true rank over 100 trials. This approach is compared against the eigengap heuristic by using ALPCAHUS with both adaptive rank schemes and report the estimated rank values from the final clustering. The eigengap approach consistently underestimated the rank regardless of the initial value, except when given the true rank value $d=6$. Our sign flipping approach provided much better rank estimates with much smaller variances between trials.
@@ V-B1 Quasar Flux Data
Quasar spectra data from the Sloan Digital Sky Survey (SDSS) Data Release 16 [39] is investigated using its DR16Q quasar catalog [40]. Each quasar has a vector of flux measurements across wavelengths that describes the intensity of observing that particular wavelength. In this dataset, the noise is heteroscedastic across the sample space (quasars) and feature space (wavelength), but for our experiments focused on a subset of data that is homoscedastic across wavelengths and heteroscedastic across quasars. The noise for each quasar is known given the measurement devices used for data collection [39]. In Fig. 7, a subset of the spectra data is shown for illustrative purposes to compare the visual differences in data quality. The preprocessing pipeline for the data included (filtering, interpolation, centering, and normalization) based on the supplementary material of [41]. Clusters are formed in these experiments by considering quasars with different properties, namely, two quasar groups ($K=2$) with different redshift values ($z_{1}=1.0-1.1,z_{2}=2.0-2.1$). Additionally, for the second group, a query for broad absorption lines (BAL) type quasars is done. Since the downloaded data is queried separately, one knows which points belong to which cluster group, meaning one can compute clustering error for comparison purposes. A training set (5000 samples total) is set aside to learn any model parameters and the rest (5000 samples total) is used for a test dataset. Trials are formed by randomly selecting 400 quasar spectra samples per group and running clustering algorithms for 50 total trials. For rank estimation, the noisy oracle approach is used to find that there is no improvement to clustering quality for rank values greater than $\hat{d}=3$, so this value is used for any applicable algorithms that require it.
@@
Figure 6(a) shows clustering error for this quasar spectra multi-cluster data. The ensemble version of ALPCAHUS was very close in clustering quality to the noisy oracle, suggesting that it accurately learned the subspace bases while clustering the data groups. The non-ensemble version of ALPCAHUS ($B=1$) had large variances in clustering error, indicating some challenges in getting close to the optimal cost function minima with this data. However, based on median values, it still managed to get a lower error than KSS, ADSSC, and TSC. Additionally, Fig. 6(b) shows the average subspace affinity error of these methods after applying LR-ALPCAH to the clusterings for all methods. ALPCAHUS better learned the subspace bases than the other methods when $B=16$, which was chosen as the smallest value that had small run-to-run variances while not taking significantly longer to compute. Both of these results show the benefit of developing heteroscedastic specific algorithms in the subspace clustering context. Table I reports median time complexity and memory requirements along with mean clustering error and mean subspace error. In this data instance, the ensemble method gets close in time to the non-ensemble methods due to our multi-threaded implementation. Yet, because of the multi-threaded implementation, the memory requirements are larger for the ensemble method as opposed to the non-ensemble method. Overall, relative to other clustering methods, ALPCAHUS is competitive time-wise at the cost of increased memory requirements.
@@ V-B2 Indian Pines Data
Hyperspectral image (HSI) segmentation data, called Indian Pines [42], is explored in this section. This image of size $145\times 145$ contains $D=200$ reflectance bands for each pixel that is around the $0.4-2.5\mu m$ wavelength. HSI data is known to be very noisy due to thermal effects, atmospheric effects, and camera electronics, leading to works that study per-pixel noise estimation in hyperspectral imaging [43]. In total, there are $K=16$ classes composed of alfalfa, corn, grass, wheat, soybean, and others. There is one additional class (background) that is withheld from clustering as commonly done [33]. Therefore, $N=10,249$ instead of $N=145\times 145$. In other works, researchers have found that $\hat{d}=5$ is an appropriate rank parameter since it covers $95\%$ of the cumulative variance [33]. A subset of the data matrix $Y\in\mathbb{R}^{200\times 10249}$ is split into $2,500$ samples to learn model parameters. From this, the learned subspaces are applied on $Y$ to cluster all data so that the heatmaps can be compared against the ground truth. The ground truth image is found in Fig. 8(a). For perspective, randomly guessing would lead to 94% clustering error due to $K=16$ clusters. For consistency, the Hungarian algorithm is again applied to the clustering results in Fig. 8 to make it easier to compare against the ground truth label image in Fig. 8(a).
@@ Methods
$K$-means is compared against in Fig. 8(b) to illustrate the difficulty of the problem. ALPCAHUS results are shown in Fig. 8(d) next to EKSS results in Fig. 8(c). Clustering error is reported for these algorithms along with mean IOU to measure the similarity between the images. ALPCAHUS achieved the lowest clustering error and highest mIOU (53% / 31%) relative to the other approaches such as $K$-means (77% / 16%) and EKSS (64% / 27%). We note that our ALPCAHUS results are similar to [33] even though, in that work, they instead treat the noise as homoscedastic and instead model the signal as heteroscedastic. Further, the EKSS results on this dataset are similar to the homoscedastic PCA approach used in [33]. This indicates some utility in modeling heteroscedasticity in hyperspectral images. Yet, the reflectance bands themselves also appear to be heteroscedastic with some bands being noisier than others [44]. Thus, developing a method that is doubly heteroscedastic with respect to both the samples and features is an interesting direction of future work.
For reference, the SOTA result is about 10% misclassification rate in a classification setting (i.e., not clustering) [45]. In a clustering setting, recent works such as [46] have achieved a 40% clustering error which is 13% lower than our results. We do not claim SOTA results with HSI data, only that heteroscedastic union of subspace modeling improved results over homoscedastic union of subspace modeling in finding reflectance band groups with similar characteristics.
@@ VI Conclusion
This paper proposed ALPCAHUS, a subspace clustering algorithm that can find subspace clusters whose samples contain heteroscedastic noise. For future work, a union of manifolds generalization could possibly be useful. For example, Alzheimer’s disease patients often have resting state functional MRI activations different than cognitively normal individuals [47], so one could consider these different classes to belong to different manifolds in the ambient space. Previous research has shown manifold learning to more correctly model temporal dynamics than subspace modeling in this domain [48]. Another direction of future work could be to consider heteroscedasticity across the feature space instead of the sample space. For example, one might be interested in biological sequencing of different species with similar genes where the gene marker counts naturally follow heteroscedastic distributions [49]. These generalizations are nontrivial so it is left for future work.
While our implementation of ALPCAHUS benefited from parallelization, it needed more computation time and memory than TSC. Because of this, there is an opportunity to further optimize our code base to remove certain matrix multiplication operations and instead use single vector dot products to reduce memory. Further, since our subspace basis step requires an SVD computation for an initial low-rank estimate, perhaps the Krylov-based Lanczos algorithm [50] can be used to further reduce time and memory for each trial leading to more significant improvements in memory usage and time. Overall, ALPCAHUS was more robust towards heteroscedastic noise than other clustering methods, making it easier to identify correct clusterings under this kind of heteroscedastic noise model.
@@ VII Acknowledgments
This work was supported in part by NSF CAREER Grant CCF-1845076 and NSF Grant CCF-2331590. There are no financial conflicts of interest in this work. We thank David Hong for suggesting heteroscedastic data applications such as astronomy spectra used in this work.
@@ VIII-A ALPCAHUS Convergence Experiment
This section focuses on providing empirical verification of the ALPCAHUS convergence theorem in Thm. 1. The quasar spectra flux data from SDSS (DR16Q catalog) is used in this experiment. The noise variance threshold parameter $\alpha$ is set to $10^{-6}$ to lower bound the noise variance estimation. Further, as shown in Thm. 1, it is necessary to use the cluster reassignment criteria in (17) that rejects repeated assignments of points to prevent cycling. For simplicity, only one base trial ($B=1$) is run in this experiment for plotting purposes and to compare against the TIPS initialized variant of ALPCAHUS. This variant is only possible when $B=1$ as discussed in the paper. In Fig. 9, the cost function value is plotted over iterations for both versions of the ALPCAHUS method.
@@
Both upper and lower bound cost function reference values are provided in Fig. 9. The upper bound is generated by first using randomly initialized subspaces as the true subspaces and secondly, using (17) to get cluster labels. From this, the cost function in (15) is computed. Likewise, a lower bound is generated by first using the known cluster labels and secondly, estimating the subspaces and noise variances with (16). As observed in Fig. 9, ALPCAHUS converges relatively quickly with only a few iterations. Note that in this figure, the stopping criteria is not implemented. In other words, ALPCAHUS is run for a fixed number of iterations. However, as clearly demonstrated, convergence to a fixed point is achieved earlier for both versions of ALPCAHUS before reaching the maximum number of iterations. This indicates the usefulness of the stopping criteria to speed up the algorithm in various applications.

### v3

@@ V-A2 Clustering Error
Table 2(h) shows the clustering quality of these algorithms for synthetic heteroscedastic data. Fig. 2 complements Table 2(h) by exploring the complete heteroscedastic landscape with a visual representation that is easier to understand. We define a method called “noisy oracle” where an oracle uses the true cluster assignments to apply PCA only on the low noise data for each cluster. Using these subspace basis estimates of each cluster, the oracle performs cluster assignments using (29). This oracle reference provides a kind of lower bound on realistic clustering performance since it can approximate subspace bases separately given the true labeling.
ALPCAHUS ($B=1$) with TIPS initialization achieved lower clustering error than KSS with TIPS initialization. For the ensemble methods ($B>1$), both EKSS and ALPCAHUS significantly improved. However, EKSS still had higher clustering error in more heteroscedastic regions. Meanwhile, ALPCAHUS remained very close to the noisy oracle, indicating it more accurately estimated the underlying true labeling. Compared to other methods like TSC and ADSSC, ALPCAHUS generally outperformed them. Overall, the ensemble version of ALPCAHUS was generally more robust against heteroscedasticity compared to other subspace clustering algorithms. To summarize, the effects of data quality and data quantity is explored on the heteroscedastic landscape. To our knowledge, this is the first systematic analysis of heteroscedasticity effects on subspace clustering quality in the literature.
@@ V-A3 Effects of Good Data
Additionally, the effects of good data quantity on the subspace clustering problem is explored to understand at what point it becomes advantageous to use ALPCAHUS over KSS methods. The following parameters are fixed $\{\nu_{1}=0.1,N_{2}=500\}$ and the following vary $\{N_{1},\nu_{2}\}$. The reporting metric is the percentage difference of EKSS clustering error (%) - ALPCAHUS clustering error (%), i.e., $\text{error}_{\text{EKSS}}-\text{error}_{\text{ALPCAHUS}}$, meaning higher values indicate that ALPCAHUS is better than EKSS. Fig. 3 shows that ALPCAHUS performed better than EKSS even up to $N_{1}=160$ which represents about 33% of the total data. ALPCAHUS never performed worse than EKSS but the advantage gap narrowed as $N_{1}$ increased. Thus, there is a wide range of conditions for which ALPCAHUS is preferable to homoscedastic methods.
@@ V-A4 Clustering Initialization
Sec. IV-D proposed using the TIPS initialization scheme in the heteroscedastic context since the dot product metric used to construct the affinity matrix in (42) is provably robust to heteroscedastic noise. We fixed the following parameters $\{N_{1},N_{2}\}$ and varied $\nu_{2}$ while keeping $\nu_{1}=0.1$. The noisy oracle result is included to establish a realistic performance baseline. Fig. 4 shows that TIPS initialization ALPCAHUS $(B=1)$ outperformed random initialization for the non-ensemble ALPCAHUS method $(B=1)$ across the entire heteroscedastic landscape. Thus, TIPS should always be used for the non-ensemble version for higher clustering accuracy. The ensemble version of ALPCAHUS $(B>1)$ is not included in these results because the ensemble process inherently takes advantage of random initialization to achieve a different labeling for each trial.
@@ V-A5 Rank Estimation of Clusters
In Sec. IV-C, a random sign flipping method is proposed to adaptively find and shrink the subspace dimension of clusters when subspace dimension is unknown. In Fig. 5, synthetic subspaces are generated with true rank $d=6$ to explore how the initial rank parameter affects the ability to estimate the true rank over 100 trials. This approach is compared against the eigengap heuristic by using ALPCAHUS with both adaptive rank schemes and report the estimated rank values from the final clustering. The eigengap approach consistently underestimated the rank regardless of the initial value, except when given the true rank value $d=6$. Our sign flipping approach provided much better rank estimates with smaller variances between trials.
@@ V-B1 Quasar Flux Data
Quasar spectra data is downloaded from the Sloan Digital Sky Survey (SDSS) Data Release 16 [sdss] using the DR16Q quasar catalog [sdss2]. Each quasar has a vector of flux measurements across wavelengths that describes the intensity of observing that particular wavelength. In this dataset, the noise is heteroscedastic across the sample space (quasars) and feature space (wavelength), but for these experiments, we focused on a subset of data that is homoscedastic across wavelengths and heteroscedastic across quasars. The noise for each quasar is known given the measurement devices used for data collection [sdss]. In Fig. 7, a subset of the spectra data is shown for illustrative purposes to compare the visual differences in data quality. The preprocessing pipeline for the data included (filtering, interpolation, centering, and normalization) based on the supplementary material of Ref. [optimal_pca]. Clusters are formed in these experiments by considering quasars with different properties, namely, two quasar groups ($K=2$) with different redshift values ($z_{1}=1.0-1.1,z_{2}=2.0-2.1$). Additionally, for the second group, only broad absorption lines (BAL) type quasar data is collected. Since the downloaded data is queried separately, clustering association is known, meaning one can compute clustering error for comparison purposes. A training set (5000 samples) is set aside to learn any model parameters and the rest (5000 samples) is used for a test dataset. Trials are formed by randomly selecting 400 quasar spectra samples per group and running clustering algorithms for 50 total trials. For rank estimation, the noisy oracle approach is used, given the known cluster labels, to find that there is no improvement to clustering quality for rank values greater than $\hat{d}=3$, so this value is used for any applicable algorithms.
@@
Fig. 6(a) shows clustering error for this quasar spectra multi-cluster data. The ensemble version of ALPCAHUS was very close in clustering quality to the noisy oracle, suggesting that it accurately learned the subspace bases while clustering the data groups. The non-ensemble version of ALPCAHUS ($B=1$) had large variances in clustering error, indicating some challenges in getting close to the optimal cost function minima with this data. However, based on median values, it still managed to get a lower error than KSS, ADSSC, and TSC. Additionally, Fig. 6(b) shows the average subspace affinity error of these methods after applying LR-ALPCAH to the clusterings for all methods. ALPCAHUS better learned the subspace bases than the other methods when $B=16$, which was chosen as the smallest value that had small run-to-run variances while not taking significantly longer to compute. Both of these results show the benefit of developing heteroscedastic-based algorithms in the subspace clustering context. Table I reports median time complexity and memory requirements along with mean clustering error and mean subspace error. In this quasar example, the ensemble method gets close in time to the non-ensemble methods due to our multi-threaded implementation. Yet, because of the multi-threaded implementation, the memory requirements are larger for the ensemble method as opposed to the non-ensemble method. Overall, relative to other clustering methods, ALPCAHUS is competitive time-wise at the cost of increased memory requirements.
@@ V-B2 Indian Pines Data
Hyperspectral image (HSI) segmentation data is explored in this section, specifically Indian Pines [pines]. This image of size $145\times 145$ contains $D=200$ reflectance bands for each pixel that is around the $0.4-2.5\mu m$ wavelength. HSI data is known to be very noisy due to thermal effects, atmospheric effects, and camera electronics, leading to works that study per-pixel noise estimation in hyperspectral imaging [hyperspectral1]. In total, there are $K=16$ classes composed of alfalfa, corn, grass, wheat, soybean, and others. There is one additional class (background) that is withheld from clustering as commonly done in other works [collas2021probabilistic]. Because of this, the data contains $N=10,249$ samples instead of $N=145\times 145$ samples. In other works, researchers have found that $\hat{d}=5$ is an appropriate rank parameter since it covers $95\%$ of the cumulative variance [collas2021probabilistic]. A subset of the data matrix $Y\in\mathbb{R}^{200\times 10249}$ is split into $2500$ samples to learn model parameters. From this, the learned subspaces are applied on $Y$ to cluster all data so that the heatmaps can be compared against the ground truth. The ground truth image is found in Fig. 8(a). For perspective, randomly guessing would lead to 94% clustering error due to $K=16$ clusters. For consistency, the Hungarian algorithm is again applied to the clustering results in Fig. 8 to make it easier to compare against the ground truth label image in Fig. 8(a).
@@ Methods
$K$-means is included in Fig. 8(b) to illustrate the difficulty of the problem. ALPCAHUS results are shown in Fig. 8(d) next to EKSS results in Fig. 8(c). Clustering error is reported for these algorithms along with mean intersection over union (IOU) to measure the similarity between the images. ALPCAHUS achieved the lowest clustering error and highest mIOU values (53%, 31%) relative to the other approaches such as $K$-means (77%, 16%) and EKSS (64%, 27%). We note that our ALPCAHUS results are similar to Ref. [collas2021probabilistic] even though, in that work, the authors treat the noise as homoscedastic and the signal itself as heteroscedastic. Further, the EKSS results on this dataset are similar to the homoscedastic PCA approach used in Ref. [collas2021probabilistic]. This indicates some utility in modeling heteroscedasticity in hyperspectral images. Yet, the reflectance bands themselves also appear to be heteroscedastic with some bands being noisier than others [hyperspectral2]. Thus, developing a method that is doubly heteroscedastic with respect to both the samples and features is an interesting direction for future work.
For reference, the state of the art result is about 10% misclassification rate in a classification setting (i.e., not clustering) [indianpines_class_sota]. In a clustering setting, recent works such as Ref. [indianpines_clustering_sota] have achieved a 40% clustering error which is 13% lower than our results. We do not claim state of the art results with this HSI data, only that heteroscedastic union of subspace modeling improved results over homoscedastic union of subspace modeling in finding reflectance band groups with similar characteristics.
@@ VI Conclusion
This paper proposed ALPCAHUS, a subspace clustering algorithm that can find subspace clusters whose samples contain heteroscedastic noise. For future work, a union of manifolds generalization by deep learning could be useful. For example, Alzheimer’s disease patients often have resting state functional MRI activations different than cognitively normal individuals [function_mri_disease], so one could consider these different classes to belong to different manifolds in the ambient space. Previous research has shown manifold learning to more correctly model temporal dynamics as opposed to subspace modeling in this domain [functional_mri_vae]. Another direction of future work could be to consider heteroscedasticity across the feature space instead of the sample space. For example, one might be interested in biological sequencing of different species with similar genes where the gene marker counts naturally follow heteroscedastic distributions [bio_hetero]. Another interesting topic would be to explore Thm. 1 in greater detail. This theorem simply states that ALPCAHUS generates a sequence of cost function values that converge; it says nothing about recovery guarantees or the quality of that solution. Thus, provably showing how the basin of attraction changes as heteroscedastic data is introduced would lead to a greater understanding of heteroscedasticity in a subspace clustering context. Additionally, per Sec. VIII-B, extending the method to handle extremely imbalanced cluster sizes would be another interesting avenue. These generalizations are nontrivial so they are left for future work.
While our implementation of ALPCAHUS benefited from parallelization, it needed more computation time and memory than TSC. Because of this, there is an opportunity to further optimize our code base to remove certain matrix multiplication operations and instead use single vector dot products to reduce memory. Further, since our subspace basis step requires an SVD computation for an initial low-rank estimate, perhaps the Krylov-based Lanczos algorithm [fast_svd] can be used to further reduce time and memory for each trial leading to more significant improvements in memory usage and time. These additions can be fruitful in big data settings. Overall, ALPCAHUS was more robust towards heteroscedastic noise than other clustering methods, making it easier to identify correct clusterings under this kind of heterogeneity model.
@@ VII Acknowledgments
This work was supported in part by the National Science Foundation (NSF) under NSF CAREER Grant CCF-1845076, NSF Grant CCF-2331590, and by the U.S. National Institutes of Health (NIH) under Grant R21 AG082204. We thank David Hong for suggesting heteroscedastic data applications such as astronomy spectra that is used in this work.
@@ VIII-A ALPCAHUS Convergence Experiment
This section focuses on providing empirical verification of the ALPCAHUS convergence theorem in Thm. 1. The quasar spectra flux data from SDSS (DR16Q catalog) is used in this experiment. The noise variance threshold parameter $\alpha$ is set to $10^{-6}$ to lower bound the noise variance estimation. Further, as shown in Thm. 1, it is necessary to use the cluster reassignment criteria in (29) that rejects repeated assignments of points to prevent cycling. For simplicity, only one base clustering ($B=1$) is used in this experiment to compare a randomly initialized ALPCAHUS trial against the TIPS initialized variant. This variant is only possible when $B=1$ as discussed in the paper. In Fig. 9, the cost function value is plotted over iterations for both versions of ALPCAHUS.
@@
Both upper and lower bound cost function reference values are provided in Fig. 9. The upper bound is generated by using randomly initialized subspaces, and by applying (29) to get cluster labels. From this, the cost function value is computed by (19). Likewise, a lower bound is generated by using the known cluster labels, and by estimating the subspaces and noise variances by (20). As observed in Fig. 9, ALPCAHUS converges relatively quickly with only a few iterations. Note that in this figure, the stopping criteria in line (10) of Alg. ALPCAHUS is purposely ignored. In other words, ALPCAHUS is run for a fixed number of iterations. However, as clearly demonstrated, convergence is achieved earlier for both versions of ALPCAHUS before reaching the maximum number of iterations. This indicates the usefulness of the stopping criteria to speed up the algorithm in various applications.
VIII-B Balanced Cluster Assumption
The ensemble extension of ALPCAHUS relies on forming an affinity matrix in (30) that is thresholded by (31) and (32). Then, spectral clustering is applied on said affinity matrix which depends on (4). This spectral clustering operation, by the normalized cut function in (4), assumes balanced cluster partitions to avoid trivial solutions such as finding a small cut that leads to a single point belonging to its own cluster and everything else being in another cluster. Thus, the proposed ALPCAHUS method may not be ideal in situations where imbalanced clusters exist.
While one could use other clustering methods for imbalanced data such as Ref. [pet_turtle] or Ref. [subspace_clustering_imbalanced], these methods are not ideal in heterogeneous settings as they do not account for heteroscedasticity in the data. Instead, we recommend using the partition cut (“Pcut”) method introduced in Ref. [spectral_clustering_imbalanced] that performs spectral clustering-like operations on a rank-modulated degree graph as a replacement for spectral clustering which uses the normalized cut function (“Ncut”) in (4). This would allow for ALPCAHUS to be more robust towards imbalanced cluster size situations. Furthermore, the non-ensemble version of ALPCAHUS in (19), may also suffer in the imbalanced cluster regime if the TIPS-based initialization scheme in Sec. IV-D is used. This is due to the spectral clustering operation in TIPS to initialize cluster labels for ALPCAHUS. Therefore, even when $B=1$, it is also recommended to apply the “Pcut” method in Ref. [spectral_clustering_imbalanced] on the affinity matrix formed by (41).
VIII-C ALPCAHUS Parameters
For Alg. ALPCAHUS, the input data matrix $Y\in\mathbb{R}^{D\times N}$ and the number of subspaces/clusters $K\in\mathbb{Z}^{+}$ are necessary inputs that are clearly data dependent. For the candidate subspace dimension $\hat{d}_{k}$, it may be the same across all clusters or vary, and it is either known via domain knowledge or can be reasonably estimated by the procedure outlined in Sec. IV-C. If it must be estimated, it is recommended to start over-parameterized rather than under-parameterized. A scree plot of $Y$ can give an indication of sufficient rank for initialization to ensure an over-parameterized state. Further, $\hat{d}_{k}$ may also be treated simply as a hyperparameter for cross validation purposes.
For the threshold parameter $q\in\mathbb{Z}^{+}$, this can be set to the subspace dimension as shown in Ref. [l2graph_subspace_clustering] and explained in Sec. IV-A, or simply treated as a hyperparameter to be optimized during cross validation; this value is dataset dependent. The number of base clusterings $B\in\mathbb{Z}^{+}$ is either fixed to $B=1$ for the non-ensemble version of ALPCAHUS, or set as high as compute time and memory allows for the ensemble version. In other words, more trials make it easier to identify the underlying relationships between similar data samples. If $B=1$, then it is recommended to use the TIPS-based initialization scheme from Sec. IV-D.
Since each subspace basis update involves LR-ALPCAH [alpcah_journal], the $T_{1}\in\mathbb{Z}^{+}$ parameter specifies the number of LR-ALPCAH iterations. For $T_{1}$, it is recommended to keep some low value, e.g., we set $T_{1}=5$ since there is no point in converging perfectly as the cluster labels will get updated in the next reassignment anyway at the $t_{2}+1$ iteration. For the non-ensemble version, it is recommended to do as many alternating updates $T_{2}\in\mathbb{Z}^{+}$ as necessary to trigger the stopping criteria in line (10) of Alg. ALPCAHUS. However, for the ensemble version of ALPCAHUS, we find that a low value such as $T_{2}=3$ works very well in practice due to the larger number of trials $B\gg 1$. This value agrees with Fig. 9 as ALPCAHUS converges in just a few iterations anyway. For more detailed and practical code-oriented function usage, refer to the documentation at github.com/javiersc1/ALPCAHUS.